OpenStudy (anonymous):

The following is the graph of y = 2x + 5.

OpenStudy (anonymous):

im almost positive this is false

OpenStudy (anonymous):

make x=0 and solve for the equation to get the y value y= 2(0)+5 y= 5 so your y value is 5 for x make y =0 0= 2x+5 -5=2x x= -5/2 and graph your answers plotting the x-axis and the y-axis

OpenStudy (anonymous):

its false?

OpenStudy (anonymous):

it is false!

OpenStudy (anonymous):

yeah its false

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y=2/1x-3 is the quation for that one